# Score Distillation via Reparametrized DDIM

### v2

@@ Appendix D ODE derivation
Appendix E Additional Intuition: Out-of-Distribution Correction
@@
Let’s consider a noisy or blurry image $\hat{x}_{0}$. Our goal is to use our updated ODE to improve it and make it look more realistic. For $t$ close to $1$, the amount of noise is so big, that for any image $\hat{x}_{0}$ its noisy version $\hat{x}_{0}+\sigma(t)\epsilon$ looks indistinguishable from the training data of the denoiser. This explains why we can successfully use diffusion models to correct out-of-distribution images.
@@
Appendix F Additional Results
F.1 Failure Cases
We provide illustrations of typical failure cases of our algorithm in Figures 16, 17, 18 and 19.
@@ F.1 Failure Cases
F.2 Additional generations
Figures 20, 21, 22 and 23 provide additional generations produced using our method.
@@ F.2 Additional generations
F.3 Additional comparison with baselines
We provide additional qualitative comparisons with other score distillation algorithms in figs. 24, 25 and 26. We compare with Dreamfusion [poole2022dreamfusion], Magic3D [lin2023magic3d], Fantasia3D [chen2023fantasia3d], Stable Score Distillation (StableSD [tang2023stable], Classifier Score Distillation (ClassifierSD [yu2023text]), Noise-Free Score Distillation (NFSD [katzir2023noise]), ProlificDreamer or Variational Score Distillation (VSD [wang2024prolificdreamer]), Interval Score Matching (ISM) [liang2023luciddreamer], and HiFA [zhu2023hifa].

### v3

@@ Appendix F ODE derivation
Appendix G Additional Intuition: Out-of-Distribution Correction
@@
Consider a noisy or blurry image $\hat{x}_{0}$. Our goal is to use our updated ODE to improve it and make it look more realistic. For $t$ close to $1$, the amount of noise is so big, that for any image $\hat{x}_{0}$ its noisy version $\hat{x}_{0}+\sigma(t)\epsilon$ looks indistinguishable from the training data of the denoiser. This explains why we can successfully use diffusion models to correct out-of-distribution images.
@@
Appendix H Additional Results
Lastly, we discuss failure cases and showcase additional generations using our methods, along with more baseline comparisons.
H.1 Failure Cases
We provide illustrations of typical failure cases of our algorithm in Figures 20, 21, 22 and 23.
@@ H.1 Failure Cases
H.2 Additional generations
Figures 24, 25, 26 and 27 provide additional generations produced using our method.
@@ H.2 Additional generations
H.3 Additional comparison with baselines
We provide additional qualitative comparisons with other score distillation algorithms in figs. 28, 29 and 30. We compare with Dreamfusion [poole2022dreamfusion], Magic3D [lin2023magic3d], Fantasia3D [chen2023fantasia3d], Stable Score Distillation (StableSD [tang2023stable], Classifier Score Distillation (ClassifierSD [yu2023text]), Noise-Free Score Distillation (NFSD [katzir2023noise]), ProlificDreamer or Variational Score Distillation (VSD [wang2024prolificdreamer]), Interval Score Matching (ISM) [liang2023luciddreamer], and HiFA [zhu2023hifa].